OpenStudy (anonymous):

I need help! Please!! "suppose that a committe has only $3000 to spend for an activity! in order to stay within buget the committe must determine the 1st day prize amount...Determine the 1st day prize amount if the committe increases the daily amount by $20 each day"

OpenStudy (nincompoop):

you can start with 20 in the first day and that could go on until 3000 is reached.

OpenStudy (anonymous):

Im sorry its given that there are only 15 days in the activity. However it is obvious that in the first day $20 are not added.

OpenStudy (nincompoop):

there you go, there's a limiter LOL

OpenStudy (anonymous):

LOL yeah

OpenStudy (nincompoop):

okay so think about this one: in 14 days you are potentially adding $280. 14 days since you said that 20 is not added on the first day. you have the sum of $3000 and you are expected to deduct $ 280 in mathematical expression that would be 20*(14) + x = 3000 use distributive property 280 + x = 3000 isolate x by subtracting 280 on both sides of the equation 280 - 280 + x = 3000 - 280 x = 3000 - 280 I think you get the idea now

OpenStudy (nincompoop):

is this clear to you? @Govani

OpenStudy (anonymous):

so just to make it clear.. It would be x=2,720 then 2,720/15= 181.3 so would the first day be $181?

OpenStudy (anonymous):

because of the 15 days given.. right?

OpenStudy (nincompoop):

(181.33 + 20) *14 + 181.33

OpenStudy (nincompoop):

you're correct

OpenStudy (nincompoop):

someone can pocket the five cents

OpenStudy (anonymous):

HAHA LOL! thanks for the help!

OpenStudy (nincompoop):

you're great, keep up the good work

OpenStudy (nincompoop):

I think this activity is just trying to get you acquainted with formulating your own mathematical expressions in order to solve problems

OpenStudy (anonymous):

Yep, Thanks!